OpenStudy (anonymous):

A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is 89m long and 66m wide. Find the area of the training field. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (pawanyadav):

Total area =area of rectangular part+ area of two semicircular part = area of rectangle+area of circle (Two semicircular=one circle)

OpenStudy (anonymous):

Could u give the image please? Thanks

OpenStudy (pawanyadav):

And diameter of circle= width of rectangular part So radius =width /2

OpenStudy (anonymous):

@Pawanyadav and @spiderghost12 I understand how to work the problem out i'm just not sure of how i need to go about finding my final answer. And sorry but for some reason it isnt letting me upload the pic :( I would just use the numbers to help you with the problem.

OpenStudy (anonymous):

@Pawanyadav You aren't making any sense to me.. sorry but I know that you take 89 and multiply it by 66 and then you you divide 66 by 2 and then multiply 3.14 times the number is gives you. Then, you will see if you need to add or subtract from there or divide that number by 2 again to get your final answer.

OpenStudy (pawanyadav):

Area =length ×width +π(width/2)^2

OpenStudy (anonymous):

Thank you so much I have a couple more to do so if you could help me with a few more after till I get it right that would be great!!

OpenStudy (pawanyadav):

Welcome

OpenStudy (anonymous):

How long will it take you to find the answer?

OpenStudy (pawanyadav):

Area=89×66 +3.14(33)^2 Calculate yourself

OpenStudy (pawanyadav):

You know rules of open study

OpenStudy (anonymous):

this isn't a test problem this is a practice problem

OpenStudy (anonymous):

thank you so much that was right could you help me do that with a couple more?

OpenStudy (anonymous):

I will fan you and give you a medal

OpenStudy (pawanyadav):

=5874+3419.46=9293.46

OpenStudy (anonymous):

A rectangular piece of paper with length 34cm and width 14cm has two semicircles cut out of it, as shown below. Find the area of the paper that remains. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

This is my new problem i am working on

OpenStudy (pawanyadav):

Same everything Just this time you have to subtract area of circle from area of rectangle.

OpenStudy (anonymous):

so 34 * 14 -3.14(7)^2?

OpenStudy (pawanyadav):

Correct

OpenStudy (anonymous):

A rectangular paperboard 29in long and 22in wide has a semicircle cut out of it, as shown below. Find the area of the paperboard that remains. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

what about this one? Add or subtract?

OpenStudy (anonymous):

Btw i just tried subtracting and adding it didnt work

OpenStudy (pawanyadav):

Subtract but this time it's not a circle It's only a semicircle. So Area=length ×width-(π11^2)/2

OpenStudy (pawanyadav):

Area of circle /2

OpenStudy (anonymous):

so what would that look like plugged in?

OpenStudy (anonymous):

I got it!!

OpenStudy (pawanyadav):

Area=29×22-π(11^2)/2 =638-189.97=448.03

OpenStudy (anonymous):

A rose garden is formed by joining a rectangle and a semicircle, as shown below. The rectangle is 23ft long and 16ft wide. Find the area of the garden. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer. This one?

OpenStudy (pawanyadav):

Previous question but add this time

OpenStudy (anonymous):

would it be 568.96? or would i divide it by 2?

OpenStudy (pawanyadav):

I need to go ...can't help more. Area=23×16+π(8^2)/2

OpenStudy (anonymous):

Okay then??? Thanks for the help anyways!!

OpenStudy (pawanyadav):

You are welcome

OpenStudy (pawanyadav):

I need to go ...can't help more. Area=23×16+π(8^2)/2

OpenStudy (pawanyadav):

Previous question but add this time

OpenStudy (anonymous):

468.48?